Trace the curve y^3 = x^2 −1 , and state all the properties you use to trace it.
1
Expert's answer
2020-03-04T17:42:02-0500

"y^3=x^2-1" ,"y=\sqrt[3]{x^2-1}"

1 thej domain "x\in(-\infty;+\infty)"

2 intersection points with coordinate axes

"Ox: y=0\\
x^2-1=0\\
x_1=1, x_2=-1\\
A(1;0), B(-1;0)\\
Oy:x=0\\
y^3=-1\\
y=-1\\
C(0;-1)"

3 symmetry

"y^3=(-x)^2-1=x^2-1"

the graph is symmetric about the axis "Oy"

4 the function is continuous

5 we find the critical points

"y'=\frac{1}{3}(x^2-1)^{-\frac{2}{3}}\cdot2x\\
x=0, x=1,x=-1\\
x\in(-\infty,-1)\cup(-1,0), y'<0, y\searrow\\
x\in(0,1)\cup(1,\infty), y'>0,y\nearrow\\"

Then "x=0" is the maximum point

"y=-1,x=0"

6 find the inflection points

"y''=\frac{2}{3}(x^2-1)^{-\frac{2}{3}}+\frac{2}{3}x\cdot(-\frac{2}{3})(x^2-1)^{-\frac{5}{3}}=\\
=\frac{2}{3}(x^2-1)^{-\frac{5}{3}}(x^2-1-\frac{4}{3}x^2)=\\
=\frac{2}{3}(x^2-1)^{-\frac{5}{3}}(-\frac{1}{3}x^2-1)\\
x\in(-\infty,-1)\cup(1,+\infty), y''<0\\
x\in (-1,1), y''>0"

7 there are no vertical asymptotes

 horizontal asymptote

"y=kx+b\\
k=\lim_{x\to\infty} \frac{\sqrt[3]{x^2-1}}{x}=0\\
b=\lim_{x\to\infty} (\sqrt[3]{x^2-1}-0\cdot x)=\infty"

there are no horizontaltical asymptotes

8 intervals for the range

"y>0, x^2-1>0\\
x^2>1\\
x\in(-\infty;-1)\cap (1;\infty)\\
y<0, x\in(-1;1)"

9 behavior at infinity

"x\to-\infty, y\to+\infty\\
x\to+\infty, y\to+\infty"